OpenStudy (anonymous):

What is the discriminant of 8x2 – 2x + 3 = 0?

OpenStudy (anonymous):

-92

OpenStudy (anonymous):

techincally an equation cant have a discriminant

OpenStudy (anonymous):

D = b^2 - 4ac = 4 - 96

OpenStudy (anonymous):

you obviously mean the discriminant of the LHS expression , and then from this this you can deduce the number of solutions to the given equation

OpenStudy (anonymous):

nevertheless, you must get your notation etc.. in order like when people say solve x^4 +4x +3 , when they mean to say "factor"